# New Approach to Strong Correlation: Twisting Hubbard into the Orbital Hatsugai-Kohmoto Model

### 16 Jan 2024 v1

@@ IV Concluding Remarks
We have shown here that orbital HK offers a powerful analytical tool for treating Mott physics described by the Hubbard model. While the agreement with Hubbard is expected in the large orbital number limit, already at $N_{\alpha}=2,4$, we find spot-on agreement with Hubbard for double occupancy at half-filling and $1/8$-hole-doping as Fig. (3c) attests. This is not a surprise. After all, we have shown that HK represents a fixed pointZhao et al. (2023b); Huang et al. (2022) for Mott physics in which not even Hubbard is a relevant perturbation. The fixed point obtains because HK contains the most relevant interaction with scaling dimension $-2$ of a Fermi liquid. The large negative scaling dimension stems from the single momentum integration. Adding orbitals to each ${\bf k}$-point does not lead to further momentum integrations as such decorations of each ${\bf k}$-point just appear as discrete summations. What the orbital content does add is non-commutativity between the kinetic and potential energy terms, thereby pushing it closer to the Hubbard model. Five manifestations of this non-commutativity stand out: 1) absence of a quasiparticle peak in the density of states at the Mott transition in agreement with standard Hubbard resultsMaier et al. (2005); Qin et al. (2016); LeBlanc et al. (2015), 2) the appearance of DSWT, 3) near-perfect quantitative agreement (see Fig. (3c)) with the QMC calculations on the Hubbard model for double occupancy that asymptotes to zero as $U$ increases, 4) a pseudogap in the density of states, and 5) separation of spin and charge excitations in the heat capacity into low and high-temperature regimes. As all of this physics is present in the Hubbard model, we conclude the orbital HK model is an effective simulator of Mott physics exemplified by the Hubbard model. This connection is apparent because the orbital extension is equivalent to replicas of finite Hubbard clusters covering the Brillouin zone all connected via TBC. Given the simplicity and cost-effective nature of our approach, orbital HK represents a new computational tool for strongly correlated Mott physics.
Acknowledgements We thank Shiwei Zhang for sending us his double occupancy data that is plotted in Fig. 3 (green dots) and George Sawatzky for a helpful e-mail exchange. This work was supported by the Center for Quantum Sensing and Quantum Materials, a DOE Energy Frontier Research Center, grant DE-SC0021238 (P. M. and P. W. P.). PWP also acknowledges NSF DMR-2111379 for partial funding of the HK work which led to these results.

### 23 Jan 2024 v2

@@ IV Concluding Remarks
We have shown here that orbital HK offers a powerful analytical and numerical tool for treating Mott physics described by the Hubbard model. While the agreement with Hubbard is expected in the large orbital number limit, already at $n=2,4$, we find spot-on agreement with Hubbard for double occupancy at half-filling and $1/8$-hole-doping as Fig. (3c) attests. This is not a surprise. After all, we have shown that HK represents a fixed pointZhao et al. (2023b); Huang et al. (2022) for Mott physics in which not even Hubbard is a relevant perturbation. The fixed point obtains because HK contains the most relevant interaction with scaling dimension $-2$ of a Fermi liquid. The large negative scaling dimension stems from the single momentum integration. Adding orbitals to each ${\bf k}$-point does not lead to further momentum integrations as such decorations of each ${\bf k}$-point just appear as discrete summations. What the orbital content does add is non-commutativity between the kinetic and potential energy terms, thereby pushing it closer to the Hubbard model. Five manifestations of this non-commutativity stand out: 1) absence of a quasiparticle peak in the density of states at the Mott transition in agreement with standard Hubbard resultsMaier et al. (2005); Qin et al. (2016); LeBlanc et al. (2015), 2) the appearance of DSWT, 3) near-perfect quantitative agreement (see Fig. (3c)) with the QMC calculations on the Hubbard model for double occupancy that asymptotes to zero as $U$ increases, 4) a pseudogap in the density of states, and 5) separation of spin and charge excitations in the heat capacity into low and high-temperature regimes. As all of this physics is present in the Hubbard model, we conclude the orbital HK model is an effective simulator of Mott physics exemplified by the Hubbard model. This connection is apparent because the orbital extension is equivalent to replicas of finite Hubbard clusters covering the Brillouin zone all connected via TBC. Given the simplicity and cost-effective nature of our approach, orbital HK represents a new computational tool for strongly correlated Mott physics.
Acknowledgements We thank Shiwei Zhang for sending us his double occupancy data that is plotted in Fig. 3d (green dots) and George Sawatzky for a helpful e-mail exchange. This work was supported by the Center for Quantum Sensing and Quantum Materials, a DOE Energy Frontier Research Center, grant DE-SC0021238 (P. M. and P. W. P.). PWP also acknowledges NSF DMR-2111379 for partial funding of the HK work which led to these results.